# Collective dynamics versus entanglement in quantum battery performance

### v2

@@ III Results
For bipartite correlations, we again compute the concurrence and the BEE. Across all interaction configurations, the concurrence exhibits a qualitatively different temporal behavior compared to the previous configuration: it starts with a finite value at $t=0$ and subsequently decreases as the instantaneous power increases [Fig. 2 ($a_{1},b_{1},c_{1}$)]. For the cases $J_{x}=1$, $J_{y}=J_{z}=0$ and $J_{x}=J_{y}=1,J_{z}=0$, the BEE grows more gradually than the instantaneous power and reaches its maximum at later times, consistent with the general trend observed in the interacting-charger setup [Fig. 2 ($a_{1},b_{1}$)]. However, when all interaction terms are present ($J_{x}=J_{y}=J_{z}=1$), the BEE remains nearly constant throughout the evolution, suggesting a suppression of entanglement dynamics despite strong coupling among the spins [Fig. 2 ($c_{1}$)]. Notably, the BEE maximum, when present, still appears after the peak in instantaneous power [Fig. 2 $(a_{1},b_{1},c_{1})$], maintaining the same temporal hierarchy as in the previous case. The early-time surge in power thus originates from rapid coherent excitation transfer between the noninteracting charger and the interacting battery, while the battery’s internal interactions primarily act to redistribute correlations locally. This results in a delayed and modest enhancement of pairwise entanglement without significantly altering the overall bipartite correlation structure of the system. However, an important qualitative difference emerges: while the instantaneous power increases, the quantum coherence in the battery decreases. This indicates that the interacting battery tends to convert quantum coherence into energetically useful, more classical population distributions during the charging process [Fig. 2 $(a_{1},b_{1},c_{1})$].
@@
We observe a strikingly similar behaviour of the instantaneous power in both configurations: a noninteracting battery charged with an interacting charger, and an interacting battery charged with a noninteracting charger (compare Fig. 1 and Fig. 2). The instantaneous power of the quantum battery is defined as $P_{i}(t)=\frac{d}{dt}\langle\hat{H}_{B}\rangle=i\langle[\hat{H},\hat{H}_{B}]\rangle$, where the total Hamiltonian is taken as $\hat{H}=\hat{H}_{B}+\hat{H}_{C}$. Since $\hat{H}_{B},\hat{H}_{B}]=0$, the power is governed entirely by the commutator $P_{i}(t)=i\langle[\hat{H}_{C},\hat{H}_{B}]\rangle$. Thus, the instantaneous power originates purely from the noncommutativity between the battery Hamiltonian and the charger Hamiltonian. In the case of a noninteracting battery with an interacting charger, the interacting nature of $\hat{H}_{C}$ ensures that $\hat{H}_{C},\hat{H}_{B}]\neq 0$, leading to a nontrivial power profile. Conversely, when the battery is interacting and the charger is noninteracting, the structure of $\hat{H}_{B}$ now plays the dominant role in producing a similar nonvanishing commutator. In both situations, the source of power is effectively the same mathematical object, the commutator between the two subsystems, resulting in qualitatively similar early-time behaviour of the instantaneous power. The differences between the two scenarios are expected to manifest only at later times, where the interacting Hamiltonian induces many-body correlations and enhanced entanglement dynamics in the system.
Figure 3: Unfair charging. (a) Parallel charging: Time evolution of the stored energy $\Delta E$ for a noninteracting battery, where each spin is charged independently by a local transverse field of strength $h_{x}=1$ and $h_{x}=N$ (with all spin-spin interactions set to zero). (b) Global charging: Time evolution of the stored energy $\Delta E$ for the same noninteracting battery, charged by an all-to-all interacting spin charger with interaction strengths $J_{x}=1$ and $J_{x}=N$, in the absence of any external transverse field. $N=8$ (open boundary conditions), with $h_{z}=1$ fixing the battery energy scale.
@@
Understanding the role of interaction structure in quantum battery charging is central to identifying genuine quantum advantages. Early theoretical studies have demonstrated that collective interactions can enhance charging power compared to parallel protocols [campaioli2017enhancing]. However, it was later clarified that such advantages may arise trivially from increasing the overall energy scale of the charger Hamiltonian rather than from intrinsically quantum effects [gyhm2022quantum]. Consequently, a fair comparison between different charging protocols requires fixing the operator norm of the charger Hamiltonian, thereby isolating the role of interaction-induced many-body dynamics.
Motivated by this perspective, we investigate how the interaction order $\kappa$, defined as the number of spins simultaneously coupled in the charger Hamiltonian, influences energy storage, instantaneous power, and quantum correlations under both unconstrained and norm-constrained (fair) charging conditions. Throughout this analysis, we consider a noninteracting battery described by Eq. (1), while the charger Hamiltonian [Eq. (2) with $J_{y}=J_{z}=0$] implements $\kappa$-body interactions.
@@
We first analyze the charging dynamics without imposing any constraint on the operator norm of the charger hamiltonian $\|\hat{H}_{C}\|$. This setting isolates classical contributions to energy storage arising purely from an increased energy scale of the charger, rather than from inherently quantum effects such as entanglement or multipartite correlations.
@@
Figure 4: Fair charging with $\bm{h_{x}\neq 0}$ and $\bm{J_{x}=h_{x}}$. Instantaneous power $P_{i}$ as a function of time $t$ for a noninteracting battery with $h_{z}=1$, charged by a $\kappa$-local interacting charger in the presence of an additional transverse field of strength $h_{x}$. The charger Hamiltonian is normalized such that $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the parameters $J_{x}$ and $h_{x}$ under the constraint $J_{x}=h_{x}$. The system size is $N=8$.
@@
Figure 5: Fair charging with $\bm{h_{x}\neq 0}$ and $\bm{J_{x}>h_{x}}$. A noninteracting battery with $h_{z}=1$ is charged by a $\kappa=1,2,3$-local interacting charger in the presence of an additional transverse field of strength $h_{x}$. The charger Hamiltonian is normalized such that $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the parameters under the constraint $J_{x}>h_{x}$. Panels (a) and (b) show, respectively, the instantaneous power $P_{i}$ and the ABEE $\overline{S}$ as functions of time $t$.
@@
To show this classical advantage, we first consider the simplest case of parallel charging, where both the battery and the charger consist of noninteracting spins. The battery Hamiltonian is given by Eq. (1) with $h_{z}=1$, and the charger Hamiltonian is defined by Eq. (2) with $J_{x}=0$, supplemented by a transverse field $h_{x}$. The operator norm of the charger scales with the applied field: for $h_{x}=h_{z}$, we have $\lVert\hat{H}_{C}\rVert=N$, while for $h_{x}=N$, it increases as $\lVert\hat{H}_{C}\rVert=N^{2}$. As a consequence, the charging strength grows with $h_{x}$. Under $h_{x}=h_{z}$, the average stored energy reaches only half of the battery’s maximum capacity, $\Delta E_{\mathrm{max}}=Nh_{z}$, because the battery Hamiltonian partially counteracts the spin flips induced by the charger [shukla2025system]. Increasing $h_{x}$ further allows the charger to dominate the dynamics, coherently flipping all spins and achieving maximal energy storage, $\Delta E_{\mathrm{max}}=2Nh_{z}$ at $h_{x}=Nh_{z}$ [Fig. 3($a$)]. Importantly, this enhancement results entirely from the classical increase in energy scale, with no contribution from quantum correlations.
@@
Next, we examine a fully interacting charger with $\kappa=N$ (all-to-all coupling) and $h_{x}=0$, while varying the interaction strength $J_{x}$. For $J_{x}=h_{z}$, the operator norm is small, $\lVert\hat{H}_{C}\rVert=1$, yielding modest maximum energy storage, $\Delta E_{\mathrm{max}}\approx 0.625$. By increasing the interaction to $J_{x}=Nh_{z}$, the operator norm rises to $\lVert\hat{H}_{C}\rVert=N$, resulting in $\Delta E_{\mathrm{max}}=Nh_{z}$, which matches the maximal stored energy achieved in the parallel scenario with $h_{x}=h_{z}$ [Fig. 3($b$)]. This comparison demonstrates that augmenting the charger strength either via the transverse field in parallel charging or via interaction strength in global charging produces equivalent increases in stored energy through classical effects alone.
@@
The equivalence of these two protocols under unconstrained norms emphasizes that the observed enhancements in the unfair charging scenario are predominantly classical. Consequently, to assess genuine quantum advantages in charging performance, it is necessary to normalize the operator norm of the charger across different configurations. This fair charging condition ensures that any observed improvement in instantaneous power or energy transfer can be attributed to collective quantum effects, such as multipartite correlations or operator spreading, rather than trivial scaling of the Hamiltonian. Under this condition, one can systematically investigate the relationship between instantaneous power and quantum correlation measures such as quantum Fisher information and average bipartite entanglement entropy to reveal truly quantum contributions to the charging process.
@@
To isolate genuine quantum effects in charging dynamics, we impose the fair charging condition by fixing the norm of the charger Hamiltonian, $\lVert\hat{H}_{C}\rVert=N$. This removes any classical advantage arising purely from differences in energy scale. Under this constraint, we study the role of the interaction order, $\kappa$, which defines the number of spins simultaneously coupled in the charger Hamiltonian, on the instantaneous power $P_{i}$.
We are fixing the charger Hamiltonina norm by considering $J_{x}=h_{x}$. For $\kappa=1$, corresponding to parallel charging of noninteracting spins, and increasing $\kappa$ to include multi-spin couplings ($\kappa=2,3,\ldots,N$), we find that $P_{i}$ initially grows, peaks, and then decreases. Importantly, with fair charging, increasing $\kappa$ alone does not increase the maximum instantaneous power beyond the parallel case [Fig. 4], indicating that multi-spin correlations by themselves are insufficient to provide a power advantage. This is consistent with the observation that classical enhancement dominates when the charger norm is unconstrained.
@@
Figure 6: Fair charging with $\bm{h_{x}=0}$. Noninteracting battery with $h_{z}=1$, charged by a $\kappa$-local interacting charger ($\kappa=1$ to NN) in the absence of a transverse field ($h_{x}=0$). The charger Hamiltonian is normalized as $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the interaction strength $J_{x}$. (a) Instantaneous power $P_{i}$, (b) QFI, and (c) ABEE $\overline{S}$ as functions of time $t$, for different participation numbers $\kappa$, as indicated in panel (c). In panel (b), the horizontal dashed lines correspond to the entanglement bounds $\kappa(x)$, with $\kappa(1),\kappa(2),\dots$ indicated in the legend.
@@
To gain insight into the underlying quantum correlations, we analyze the QFI and the ABEE for all considered $\kappa$, while keeping $\lVert\hat{H}_{C}\rVert=N$ fixed. Both quantities increase with the interaction range and exhibit similar qualitative behavior [Figs. 6(b) and 6(c)], reflecting the growing multipartite nature of the quantum state as interactions become more global. Importantly, the peaks of the QFI and ABEE occur well after the instantaneous power reaches its maximum. This temporal mismatch indicates that multipartite entanglement is not the primary driver of peak power generation. Instead, collective interactions enable rapid energy transfer first, with entanglement building up subsequently as a consequence of the many-body dynamics.
@@
Figure 7: Fair charging with increasing participation number: Noninteracting battery with $h_{z}=1$, charged by a 22-local interacting charger in the absence of a transverse field ($h_{x}=0$). The charger Hamiltonian is normalized as $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the interaction strengths $J_{x_{1}}$ and $J_{x_{2}}$, corresponding to NN and NNN couplings, respectively, with $J_{x_{1}}>J_{x_{2}}$.
We now examine how extending the interaction range (pariticipation number) in the charger affects the charging dynamics under fair charging conditions. This is achieved by incorporating next-nearest-neighbor (NNN) couplings in the NN interaction, which increase the number of spins participating in the interaction without introducing fully global connectivity. The interaction part of the charging Hamiltonian is modified as
@@
We compare the instantaneous power of chargers with NN interactions alone to those including both NN and NNN couplings, maintaining a fixed operator norm, $\lVert H_{C}\rVert=N$. The results indicate that adding NNN interactions decreases the maxima of instantaneous power $P_{i}$ compared to the NN-only protocol [Fig. 7]. This demonstrates that increasing the participation number through short-range interaction extensions does not automatically translate into enhanced charging power.

### v3

@@ III Results
For bipartite correlations, we again compute the concurrence and the BEE. Across all interaction configurations, the concurrence exhibits a qualitatively different temporal behavior compared to the previous configuration: it starts with a finite value at $t=0$ and subsequently decreases as the instantaneous power increases [Fig. 3 ($a_{1},b_{1},c_{1}$)]. For the cases $J_{x}=1$, $J_{y}=J_{z}=0$ and $J_{x}=J_{y}=1,J_{z}=0$, the BEE grows more gradually than the instantaneous power and reaches its maximum at later times, consistent with the general trend observed in the interacting-chargoid setup [Fig. 3 ($a_{1},b_{1}$)]. However, when all interaction terms are present ($J_{x}=J_{y}=J_{z}=1$), the BEE remains nearly constant throughout the evolution, suggesting a suppression of entanglement dynamics despite strong coupling among the spins [Fig. 3 ($c_{1}$)]. Notably, the BEE maximum, when present, still appears after the peak in instantaneous power [Fig. 3 $(a_{1},b_{1},c_{1})$], maintaining the same temporal hierarchy as in the previous case. The early-time surge in power thus originates from rapid coherent excitation transfer between the noninteracting chargoid and the interacting battery, while the battery’s internal interactions primarily act to redistribute correlations locally. This results in a delayed and modest enhancement of pairwise entanglement without significantly altering the overall bipartite correlation structure of the system. However, an important qualitative difference emerges: while the instantaneous power increases, the quantum coherence in the battery decreases. This indicates that the interacting battery tends to convert quantum coherence into energetically useful, more classical population distributions during the charging process [Fig. 3 $(a_{1},b_{1},c_{1})$].
Figure 3: Interacting battery charged by noninteracting chargoid. Bipartite entanglement $(a_{1},b_{1},c_{1})$: Time evolution of the instantaneous power $P_{i}$, concurrence between the first and last spins $C(1,N)$, and BEE $S_{N/2}$. Tripartite entanglement $(a_{2},b_{2},c_{2})$: Time evolution of $P_{i}$ and TMI $I_{3}$. Multipartite entanglement $(a_{3},b_{3},c_{3})$: Time evolution of $P_{i}$, QFI $F_{Q}$, and ABEE $\overline{S}_{N/2}$. Panels correspond to: $(a_{1},a_{2},a_{3})$ $J_{x}=1$, $J_{y}=J_{z}=0$; $(b_{1},b_{2},b_{3})$ $J_{x}=J_{y}=1$, $J_{z}=0$; $(c_{1},c_{2},c_{3})$ $J_{x}=J_{y}=J_{z}=1$. Number of spins: $N=8$ and $h_{x}=h_{z}=1$.
@@
We observe a strikingly similar behaviour of the instantaneous power in both configurations: a noninteracting battery charged with an interacting chargoid, and an interacting battery charged with a noninteracting chargoid (compare Fig. 2 and Fig. 3). The instantaneous power of the quantum battery is defined as $P_{i}(t)=\frac{d}{dt}\langle\hat{H}_{B}\rangle=i\langle[\hat{H},\hat{H}_{B}]\rangle$, where the total Hamiltonian is taken as $\hat{H}=\hat{H}_{B}+\hat{H}_{C}$. Since $\hat{H}_{B},\hat{H}_{B}]=0$, the power is governed entirely by the commutator $P_{i}(t)=i\langle[\hat{H}_{C},\hat{H}_{B}]\rangle$. Thus, the instantaneous power originates purely from the noncommutativity between the battery Hamiltonian and the chargoid Hamiltonian. In the case of a noninteracting battery with an interacting chargoid, the interacting nature of $\hat{H}_{C}$ ensures that $\hat{H}_{C},\hat{H}_{B}]\neq 0$, leading to a nontrivial power profile. Conversely, when the battery is interacting and the chargoid is noninteracting, the structure of $\hat{H}_{B}$ now plays the dominant role in producing a similar nonvanishing commutator. In both situations, the source of power is effectively the same mathematical object, the commutator between the two subsystems, resulting in qualitatively similar early-time behaviour of the instantaneous power. The differences between the two scenarios are expected to manifest only at later times, where the interacting Hamiltonian induces many-body correlations and enhanced entanglement dynamics in the system.
Figure 4: Unfair charging. (a) Parallel charging: Time evolution of the stored energy $\Delta E$ for a noninteracting battery, where each spin is charged independently by a local transverse field of strength $h_{x}=1$ and $h_{x}=N$ (with all spin-spin interactions set to zero). (b) Global charging: Time evolution of the stored energy $\Delta E$ for the same noninteracting battery, charged by an all-to-all interacting spin chargoid with interaction strengths $J_{x}=1$ and $J_{x}=N$, in the absence of any external transverse field. $N=8$ (open boundary conditions), with $h_{z}=1$ fixing the battery energy scale.
@@
Understanding the role of interaction structure in quantum battery charging is central to identifying genuine quantum advantages. Early theoretical studies have demonstrated that collective interactions can enhance charging power compared to parallel protocols [campaioli2017enhancing]. However, it was later clarified that such advantages may arise trivially from increasing the overall energy scale of the chargoid Hamiltonian rather than from intrinsically quantum effects [gyhm2022quantum]. Consequently, a fair comparison between different charging protocols requires fixing the operator norm of the chargoid Hamiltonian, thereby isolating the role of interaction-induced many-body dynamics.
Motivated by this perspective, we investigate how the interaction order $\kappa$, defined as the number of spins simultaneously coupled in the chargoid Hamiltonian, influences energy storage, instantaneous power, and quantum correlations under both unconstrained and norm-constrained (fair) charging conditions. Throughout this analysis, we consider a noninteracting battery described by Eq. (1), while the chargoid Hamiltonian [Eq. (2) with $J_{y}=J_{z}=0$] implements $\kappa$-body interactions.
@@
We first analyze the charging dynamics without imposing any constraint on the operator norm of the chargoid Hamiltonian $\|\hat{H}_{C}\|$. This setting isolates classical contributions to energy storage arising purely from an increased energy scale of the chargoid, rather than from inherently quantum effects such as entanglement or multipartite correlations.
@@
Figure 5: Fair charging with $\bm{h_{x}\neq 0}$ and $\bm{J_{x}=h_{x}}$. Instantaneous power $P_{i}$ as a function of time $t$ for a noninteracting battery with $h_{z}=1$, charged by a $\kappa$-local interacting chargoid in the presence of an additional transverse field of strength $h_{x}$. The chargoid Hamiltonian is normalized such that $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the parameters $J_{x}$ and $h_{x}$ under the constraint $J_{x}=h_{x}$. The system size is $N=8$.
@@
Figure 6: Fair charging with $\bm{h_{x}\neq 0}$ and $\bm{J_{x}>h_{x}}$. A noninteracting battery with $h_{z}=1$ is charged by a $\kappa=1,2,3$-local interacting chargoid in the presence of an additional transverse field of strength $h_{x}$. The chargoid Hamiltonian is normalized such that $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the parameters under the constraint $J_{x}>h_{x}$. Panels (a) and (b) show, respectively, the instantaneous power $P_{i}$ and the ABEE $\overline{S}$ as functions of time $t$.
@@
To show this classical advantage, we first consider the simplest case of parallel charging, where both the battery and the chargoid consist of noninteracting spins. The battery Hamiltonian is given by Eq. (1) with $h_{z}=1$, and the chargoid Hamiltonian is defined by Eq. (2) with $J_{x}=0$, supplemented by a transverse field $h_{x}$. The operator norm of the chargoid scales with the applied field: for $h_{x}=h_{z}$, we have $\lVert\hat{H}_{C}\rVert=N$, while for $h_{x}=N$, it increases as $\lVert\hat{H}_{C}\rVert=N^{2}$. As a consequence, the charging strength grows with $h_{x}$. Under $h_{x}=h_{z}$, the average stored energy reaches only half of the battery’s maximum capacity, $\Delta E_{\mathrm{max}}=Nh_{z}$, because the battery Hamiltonian partially counteracts the spin flips induced by the chargoid [shukla2025optimizing]. Increasing $h_{x}$ further allows the chargoid to dominate the dynamics, coherently flipping all spins and achieving maximal energy storage, $\Delta E_{\mathrm{max}}=2Nh_{z}$ at $h_{x}=Nh_{z}$ [Fig. 4($a$)]. Importantly, this enhancement results entirely from the classical increase in energy scale, with no contribution from quantum correlations.
@@
Next, we examine a fully interacting chargoid with $\kappa=N$ (all-to-all coupling) and $h_{x}=0$, while varying the interaction strength $J_{x}$. For $J_{x}=h_{z}$, the operator norm is small, $\lVert\hat{H}_{C}\rVert=1$, yielding modest maximum energy storage, $\Delta E_{\mathrm{max}}\approx 0.625$. By increasing the interaction to $J_{x}=Nh_{z}$, the operator norm rises to $\lVert\hat{H}_{C}\rVert=N$, resulting in $\Delta E_{\mathrm{max}}=Nh_{z}$, which matches the maximal stored energy achieved in the parallel scenario with $h_{x}=h_{z}$ [Fig. 4($b$)]. This comparison demonstrates that augmenting the chargoid strength either via the transverse field in parallel charging or via interaction strength in global charging produces equivalent increases in stored energy through classical effects alone.
@@
The equivalence of these two protocols under unconstrained norms emphasizes that the observed enhancements in the unfair charging scenario are predominantly classical. Consequently, to assess genuine quantum advantages in charging performance, it is necessary to normalize the operator norm of the chargoid across different configurations. This fair charging condition ensures that any observed improvement in instantaneous power or energy transfer can be attributed to collective quantum effects, such as multipartite correlations or operator spreading, rather than trivial scaling of the Hamiltonian. Under this condition, one can systematically investigate the relationship between instantaneous power and quantum correlation measures such as quantum Fisher information and average bipartite entanglement entropy to reveal truly quantum contributions to the charging process.
@@
To isolate genuine quantum effects in charging dynamics, we impose the fair charging condition by fixing the norm of the chargoid Hamiltonian, $\lVert\hat{H}_{C}\rVert=N$. This removes any classical advantage arising purely from differences in energy scale. Under this constraint, we study the role of the interaction order, $\kappa$, which defines the number of spins simultaneously coupled in the chargoid Hamiltonian, on the instantaneous power $P_{i}$.
We are fixing the chargoid Hamiltonina norm by considering $J_{x}=h_{x}$. For $\kappa=1$, corresponding to parallel charging of noninteracting spins, and increasing $\kappa$ to include multi-spin couplings ($\kappa=2,3,\ldots,N$), we find that $P_{i}$ initially grows, peaks, and then decreases. Importantly, with fair charging, increasing $\kappa$ alone does not increase the maximum instantaneous power beyond the parallel case [Fig. 5], indicating that multi-spin correlations by themselves are insufficient to provide a power advantage. This is consistent with the observation that classical enhancement dominates when the chargoid norm is unconstrained.
@@
Figure 7: Fair charging with $\bm{h_{x}=0}$. Noninteracting battery with $h_{z}=1$, charged by a $\kappa$-local interacting chargoid ($\kappa=1$ to NN) in the absence of a transverse field ($h_{x}=0$). The chargoid Hamiltonian is normalized as $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the interaction strength $J_{x}$. (a) Instantaneous power $P_{i}$, (b) QFI, and (c) ABEE $\overline{S}$ as functions of time $t$, for different participation numbers $\kappa$, as indicated in panel (c). In panel (b), the horizontal dashed lines correspond to the entanglement bounds $\kappa(x)$, with $\kappa(1),\kappa(2),\dots$ indicated in the legend.
@@
To gain insight into the underlying quantum correlations, we analyze the QFI and the ABEE for all considered $\kappa$, while keeping $\lVert\hat{H}_{C}\rVert=N$ fixed. Both quantities increase with the interaction range and exhibit similar qualitative behavior [Figs. 7(b) and 7(c)], reflecting the growing multipartite nature of the quantum state as interactions become more global. Importantly, the peaks of the QFI and ABEE occur well after the instantaneous power reaches its maximum. This temporal mismatch indicates that multipartite entanglement is not the primary driver of peak power generation. Instead, collective interactions enable rapid energy transfer first, with entanglement building up subsequently as a consequence of the many-body dynamics.
@@
Figure 8: Fair charging with increasing participation number: Noninteracting battery with $h_{z}=1$, charged by a 22-local interacting chargoid in the absence of a transverse field ($h_{x}=0$). The chargoid Hamiltonian is normalized as $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the interaction strengths $J_{x_{1}}$ and $J_{x_{2}}$, corresponding to NN and NNN couplings, respectively, with $J_{x_{1}}>J_{x_{2}}$.
We now examine how extending the interaction range (pariticipation number) in the chargoid affects the charging dynamics under fair charging conditions. This is achieved by incorporating next-nearest-neighbor (NNN) couplings in the NN interaction, which increase the number of spins participating in the interaction without introducing fully global connectivity. The interaction part of the charging Hamiltonian is modified as
@@
We compare the instantaneous power of chargoids with NN interactions alone to those including both NN and NNN couplings, maintaining a fixed operator norm, $\lVert H_{C}\rVert=N$. The results indicate that adding NNN interactions decreases the maxima of instantaneous power $P_{i}$ compared to the NN-only protocol [Fig. 8]. This demonstrates that increasing the participation number through short-range interaction extensions does not automatically translate into enhanced charging power.